Peterson Company estimates that overhead costs for the next year will be $3,500,000 for indirect labor and $940,000 for factory utilities. The company uses machine hours as its overhead allocation base. If 100,000 machine hours are planned for this next year, what is the company's plantwide overhead rate? (Round your answer to two decimal places.)

Multiple Choice

  • $0.03 per machine hour.

  • $44.40 per machine hour.

  • $34.62 per machine hour.

  • $9.40 per machine hour.

  • $0.11 per machine hour.
Answer #1

Total factory overhead=3500000+940000 =4440000

Plantwide overhead rate =Total factory overhead/machine hours =4440000/100000 =44.40 per machine hour